OpenStudy (anonymous):

an isosceles triangle has a base, b and two equal legs, l. The formula below will find the perimeter of the triangle. P = 2L+B Find the length of the base if the length of each leg is 12cm and the perimeter is 31cm

OpenStudy (anonymous):

7

OpenStudy (anonymous):

2*12+x=31 24+x=31 cancel the 24 x=31-24=7 x=7